# Neural-Symbolic Collaborative Distillation: Advancing Small Language Models for Complex Reasoning Tasks

### v1

@@ Abstract
In this paper, we propose Neural-Symbolic Collaborative Distillation (NesyCD), a novel knowledge distillation method for learning the complex reasoning abilities of Large Language Models (LLMs, e.g., >13B). We argue that complex reasoning tasks are difficult for Small Language Models (SLMs, e.g., $\leq$ 7B), as these tasks demand not only general cognitive abilities but also specialized knowledge, which is often sparse and difficult for these neural-based SLMs to effectively capture. Therefore, NesyCD distills the general capabilities and specialized knowledge in LLMs using different manners. On the one hand, we distill only general abilities from teacher LLMs into the student SLMs of parameterized neural networks. On the other hand, for the specialized abilities and uncommon knowledge of a complex reasoning task, we employ a symbolic knowledge distillation approach to obtain and store the specialized knowledge within a symbolic knowledge base (KB). By decoupling general and specialized capabilities, the proposed NesyCD can achieve superior performance cost-effectively, utilizing smaller models and blending parameterized neural networks with symbolic KB. Moreover, the specialized KB generalizes well and is comprehended and manipulated by humans. Our experiments show that NesyCD significantly boosts SLMs’ complex reasoning performance on in-domain (BBH, GSM8K) and out-of-domain (AGIEval, ARC) datasets. Notably, our approach enabled the LLaMA3-8B and Qwen2-7B to surpass GPT-3.5-turbo in performance and come close to matching LLaMA3-70B, despite the latter having nine times more parameters. Our code will be available at https://github.com/Xnhyacinth/NesyCD.
@@ 1 Introduction
Large Language Models (LLMs) (Yang et al. 2024) excel in various complex reasoning tasks such as mathematical (Xia et al. 2024), commonsense (Zhao, Lee, and Hsu 2023) and symbolic reasoning (Dave et al. 2024) with In-Context Learning (ICL) (Ye et al. 2023; Shum, Diao, and Zhang 2023) and Chain-of-Thought (CoT) Prompting (Kaplan et al. 2020). Due to the high computational costs and expensive API calls required for LLMs (e.g., ChatGPT and LLaMA3-70B), enhancing Small Language Models (SLMs, e.g., $\leq$ 7B) to handle complex reasoning efficiently is more practical and crucial for large-scale deployment.
@@
  • We develop a bespoke learning strategy that incorporates novel auxiliary tasks for multi-task learning, alongside tailored error analysis augmented data techniques.
@@ 4.5 Efficiency on Dataset and Model Sizes
Regarding model size efficiency, as shown at the top of Figure 3, NesyCD outperforms other baselines across various model scales. Notably, NesyCD enables Qwen2-7B to surpass the teacher GPT-3.5 Turbo in both ID and OOD performance, despite having over 10×10$\times$10 × fewer parameters. While larger models generally show improved performance, reducing differences between methods, NesyCD provides a substantial advantage for smaller models like 1.5B and 0.5B. These results highlight NesyCD’s significant practical benefits in resource-constrained environments, as it reduces the computational cost required for SLMs while achieving performance levels that exceed those of larger LLMs. This further illustrates that SLMs (0.5B) cannot fully utilize the CoT reasoning generated by LLMs, thereby necessitating the implementation of our proposed NesyCD.
@@ 5 Conclusion
In this work, we introduce Neural-Symbolic Collaborative Distillation (NesyCD), a method aimed at enhancing the capabilities of Small Language Models (SLMs) for complex reasoning tasks that require additional knowledge and advanced reasoning skills. NesyCD leverages Large Language Models (LLMs) to analyze errors made by SLMs and generate specialized knowledge, including learning summaries and supplementary knowledge, which are organized into an external knowledge base (KB). By integrating parameter updates with the retrieval of external specialized knowledge, NesyCD improves both rationale generation and answer accuracy for SLMs. Empirical experiments show that NesyCD surpasses fine-tuning and CoT distillation baselines, in both in-domain and out-of-domain scenarios. The external KB helps identify and address gaps in SLMs’ reasoning capabilities and knowledge, offering a path for advancing SLM performance in complex reasoning tasks requiring comprehensive domain-specific understanding.

### v2

@@ Abstract
In this paper, we propose Neural-Symbolic Collaborative Distillation (NesyCD), a novel knowledge distillation method for learning the complex reasoning abilities of Large Language Models (LLMs, e.g., >13B). We argue that complex reasoning tasks are difficult for Small Language Models (SLMs, e.g., $\leq$ 7B), as these tasks demand not only general cognitive abilities but also specialized knowledge, which is often sparse and difficult for these neural-based SLMs to effectively capture. Therefore, NesyCD distills the general capabilities and specialized knowledge in LLMs in different ways. On the one hand, we distill only general abilities from teacher LLMs into the student SLMs of parameterized neural networks. On the other hand, for the specialized abilities and uncommon knowledge of a complex reasoning task, we employ a symbolic knowledge distillation approach to obtain and store the specialized knowledge within a symbolic knowledge base (KB). By decoupling general and specialized capabilities, the proposed NesyCD can achieve superior performance cost-effectively, utilizing smaller models and blending parameterized neural networks with symbolic KB. Moreover, the specialized KB generalizes well and is comprehended and manipulated by humans. Our experiments show that NesyCD significantly boosts SLMs’ complex reasoning performance on in-domain (BBH, GSM8K) and out-of-domain (AGIEval, ARC) datasets. Notably, our approach enabled the LLaMA3-8B and Qwen2-7B to surpass GPT-3.5-turbo in performance and come close to matching LLaMA3-70B, despite the latter having 9×9$\times$9 × more parameters. Our code will be available at https://github.com/Xnhyacinth/NesyCD.
@@ 1 Introduction
Large Language Models (LLMs) (Yang et al. 2024) excel in various complex reasoning tasks such as mathematical (Yu et al. 2023a), commonsense (Zhao, Lee, and Hsu 2023) and symbolic reasoning (Suzgun et al. 2022) with In-Context Learning (ICL) (Ye et al. 2023) and Chain-of-Thought (CoT) Prompting (Kaplan et al. 2020). Due to the high computational costs and expensive API calls required for LLMs (e.g., ChatGPT and LLaMA3-70B), enhancing Small Language Models (SLMs, e.g., $\leq$ 7B) to handle complex reasoning efficiently is more practical and crucial for large-scale deployment.
@@ 4.5 Efficiency on Dataset and Model Sizes
Regarding model size efficiency, as shown at the top of Figure 3, NesyCD outperforms other baselines across various model scales. Notably, NesyCD enables Qwen2-7B to surpass the teacher GPT-3.5 Turbo in both ID and OOD performance, despite having over 10×10$\times$10 × fewer parameters. These results highlight NesyCD’s significant practical benefits in resource-constrained environments, as it reduces the computational cost required for SLMs while achieving performance levels that exceed those of larger LLMs. This further illustrates that SLMs cannot fully utilize the CoT reasoning generated by LLMs, thereby necessitating the implementation of our proposed NesyCD.
@@ 5 Conclusion
In this work, we introduce Neural-Symbolic Collaborative Distillation (NesyCD), a method aimed at enhancing the capabilities of Small Language Models (SLMs) for complex reasoning tasks that require additional knowledge and advanced reasoning skills. NesyCD leverages Large Language Models (LLMs) to analyze errors made by SLMs and generate specialized knowledge, including learning summaries and supplementary knowledge, which are organized into an external knowledge base (KB). By integrating parameter updates with the retrieval of external specialized knowledge, NesyCD improves both rationale generation and answer accuracy for SLMs. Empirical experiments show that NesyCD surpasses fine-tuning and CoT distillation baselines, in both in-domain and out-of-domain scenarios.
Acknowledgements
This work was supported by the National Key R&D Program of China (No. 2022ZD0160503) and the National Natural Science Foundation of China (No.62376270, No.62276264).